# Are penalty shootouts better than a coin toss? Evidence from European football

### v1

@@ Abstract
Penalty shootouts play an important role in the knockout stage of major football tournaments, especially since the 2021/22 season, when the Union of European Football Associations (UEFA) scrapped the away goals rule in its club competitions. Inspired by this rule change, our paper examines whether the outcome of a penalty shootout can be predicted in UEFA club competitions. Based on all shootouts between 2000 and 2025, we find no evidence for the effect of the kicking order, the field of the match, and psychological momentum. In contrast to previous results, stronger teams, defined first by Elo ratings, do not perform better than their weaker opponents. Consequently, penalty shootouts are equivalent to a perfect lottery in top European football.
@@
Keywords: Elo rating; football; home advantage; momentum; penalty shootout
@@ 1 Introduction
The penalty shootout is the ultimate tie-breaking rule in (association) football and several other sports to decide which team qualifies in a knockout match. Penalty shootouts are usually kicked if the result is tied in both regular and extra time. Penalty shootouts directly follow regular time in some tournaments, too. For instance, in the 2025 Leagues Cup, games went straight to a penalty shootout after a tie, and teams were awarded three points for a win in regular time, two points for a win by a penalty shootout, and one point for a loss by a penalty shootout (Leagues Cup,, 2025).
@@
All of these issues have partially been investigated in the existing literature, as revealed by Section 2. However, our research has three innovative aspects. First, we analyse a relatively homogeneous sample provided by matches played in UEFA club competitions without any matches from national cups, where the strengths of the opponents can vary widely. Second, the comeback team is identified in a novel way, according to which team scored the last goal. Third, team strength is measured by Football Club Elo Ratings, a well-established statistical method to create ratings based on past performance (Aldous,, 2017; Gomes de Pinho Zanco et al.,, 2024; van Eetvelde and Ley,, 2019). Previous studies used either the division of the team (Arrondel et al.,, 2019; Krumer,, 2020), or betting odds (Wunderlich et al.,, 2020; Pipke,, 2025) for this purpose. But the first approach is inappropriate in international competitions, while betting odds suffer from several biases (Braun and Kvasnicka,, 2013; Cain et al.,, 2000; Feddersen et al.,, 2017; Winkelmann et al.,, 2021), and may be less persuasive for decision-makers than a measure given by an “objective” mathematical formula.
As we show, the outcomes of the 268 penalty shootouts played in UEFA club competitions between 2000 and 2025 seem to be entirely unpredictable. This contradicts the results of some previous works, which uncover a first-mover advantage (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020) and the positive impact of psychological momentum (Krumer,, 2021). Furthermore, according to our knowledge, all existing studies agree that stronger teams are more likely to win in a shootout (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025).
@@ 2 Related literature
The issue of first-mover advantage in football penalty shootouts has received serious attention since the pioneering work of Apesteguia and Palacios-Huerta, (2010). Some studies find evidence for a significant advantage enjoyed by the team that kicks the first penalty in each round (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020). These results have inspired formal modelling of the first-mover advantage, as well as several proposals for alternative mechanisms to mitigate or even eliminate this source of unfairness (Palacios-Huerta,, 2012; Echenique,, 2017; Brams and Ismail,, 2018; Vandebroek et al.,, 2018; Del Giudice,, 2019; Anbarcı et al.,, 2021; Csató, 2021a, ; Csató, 2021b, ; Csató and Petróczy,, 2022; Brams et al.,, 2024). The alternating (or ABBA) order was even tried in various tournaments between 2017 and 2018 (Palacios-Huerta,, 2020; Csató, 2021a, ), albeit the IFAB (International Football Association Board), the rule-making body of football, stopped the experiments in 2018 (FIFA,, 2018).
On the other hand, several papers report no significant difference between the winning probability of the first-mover and the second-mover (Kocher et al.,, 2012; Arrondel et al.,, 2019; Santos,, 2023). Kassis et al., (2021) conclude that teams whose captains win the coin toss and can choose the shooting order enjoy the advantage. The conflicting results can probably be attributed to small sample sizes (Vandebroek et al.,, 2018). Two recent studies (Vollmer et al.,, 2024; Pipke,, 2025) analysing the highest number of penalty shootouts with respect to first-mover advantage—1759 and 7116, respectively—do not find evidence for the effect of the shooting sequence. However, while the dataset of Pipke, (2025) includes all available shootouts up to the 2023/24 season, this does not mean that first-mover advantage cannot emerge in a more homogeneous subset of matches.
@@
We know only one study on the impact of psychological momentum. Krumer, (2021) analyses 214 penalty shootouts from two-legged matches in European cups (European Champion Clubs’ Cup/UEFA Champions League, European Cup Winners’ Cup/UEFA Cup Winners’ Cup, UEFA Cup/UEFA Europa League) between 1970 and 2018. Every additional goal scored by the away team in regular time in the second leg increases its probability of winning the penalty shootout by 12.5 percentage points. However, if the home team wins in regular time, it has only a 50% chance to qualify after a shootout.
@@
Stronger football teams have a significantly higher chance to win a penalty shootout based on 1067 shootouts from the domestic cup competitions of ten European associations (the top five, as well as Belgium, Portugal, Russia, Turkey, Ukraine), the Brazilian cup, the UEFA Champions League, and the UEFA Europa League (Wunderlich et al.,, 2020). However, the effect of team strength on success remains rather small; the winning probability does not exceed 60% even against an extremely weak team. The novelty of Wunderlich et al., (2020) resides in the measure of strength difference, determined by winning probabilities based on pre-match betting odds, adjusted for home advantage. The recent analysis of all available penalty shootouts until the 2023/24 season reinforces that the favourites identified by pre-match odds are more likely to win (Pipke,, 2025, Table A.5). But the favourite team might not be the stronger team due to home advantage (Wunderlich et al.,, 2020).
@@ 3 Data and methodology
Champions League Europa League Conference League Total Qualification 41 103 63 207 Knockout stage 14 27 9 50 Final 6 5 0 11 Total 61 135 72 268 UEFA Conference League was called UEFA Europa Conference League until the 2023/24 season. Knockout stage matches contain only two-legged matches, but not the final, which is played on a neutral field.
@@
Figure 1 presents the number of penalty shootouts in each season from 2000/01 to 2025/26 (in the latter case, without the spring of 2026). Besides the natural fluctuation, a dramatic increase can be seen from 2020/21, when most qualification matches were played in one leg instead of two due to COVID-19 restrictions such that the field of the match was decided by a random draw. In addition, the away goals rule was abolished, and a new competition, the UEFA (Europa) Conference League, started in 2021/22. Therefore, our sample is dominated by recent matches over these 25 years: 139 of the 268 penalty shootouts (51.9%) took place in the last six seasons.
@@
The first-mover advantage can be analysed on the basis of all (268) penalty shootouts since this information is always available. On the other hand, some matches were played on a neutral field. In order to examine home advantage, we consider two sets of matches, an extended and a restricted set, where the latter is a strict subset of the former. The extended set does not contain the finals except for the 2011/12 UEFA Champions League final, which was played in München by Bayern München and Chelsea (10 matches are dropped). In addition, one match played by an Israeli and two matches played by Belarussian teams in 2024 and 2025 are disregarded, which leads to 255 matches in the extended set. In the restricted set, 20 matches played behind closed doors in the 2020/21 season, as well as the 2011/12 UEFA Champions League final, are removed.
We identify a comeback team in two different ways. The first adopts the method of Krumer, (2021): the team that scored more goals than its opponent in the match is the comeback team. Consequently, 121 matches where the second leg game is tied should be ignored. The second approach is novel: the team that scored the last equalising goal, which is responsible for reaching the penalty shootout, is the comeback team. Hence, only 37 shootouts could not be examined.
@@ 4 Results and discussion
Table 2: Logistic regression models: the effect of and team strength
Strength difference $\Delta$ Elo rating ($\Delta^{(1)}$) Winning probability ($\Delta^{(2)}$) Coefficient -0.121 -0.027 $\beta_{0}$ Standard error 0.123 0.307 $p$-value 0.326 0.930 Coefficient -0.004 -0.194 $\beta_{1}$ Standard error 0.069 0.582 $p$-value 0.959 0.739 Number of observations 266 266 In the model Elo rating, strength difference $\Delta^{(1)}$ is the Elo rating of the home team minus the Elo rating of the away team divided by 100. In the model Elo rating, strength difference $\Delta^{(2)}$ is the winning probability of the home team based on the Elo ratings of the two teams. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
The lack of positive impact is in stark contrast to the findings of the previous literature: all existing papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025) demonstrate a significantly higher winning probability for a “sufficiently” stronger team. Thus, Table 2 investigates the same issue by logistic regressions similar to Wunderlich et al., (2020). Both coefficients are insignificant, that is, neither the field of the match nor the strength of the teams influence success in penalty shootouts, independent of how strength difference is measured. Furthermore, the estimations of $\beta_{0}$ and $\beta_{1}$ are negative in accordance with Figures 3 and 4, making the implications regarding home advantage and team strength even more robust.
@@ A.2 Supplementary tables
Table A.1: Results of binomial tests: first-mover advantage, home advantage, momentum

### v2

@@ Abstract
Penalty shootouts play an important role in the knockout stage of major football tournaments, especially since the 2021/22 season, when the Union of European Football Associations (UEFA) scrapped the away goals rule in its club competitions. Inspired by this rule change, our paper examines whether the outcome of a penalty shootout can be predicted in UEFA club competitions. Based on all shootouts between 2000 and 2025, we find no evidence for the effect of the kicking order, the field of the match, or psychological momentum. In contrast to previous results, stronger teams, defined first by Elo ratings, do not perform better than their weaker opponents. Consequently, penalty shootouts are equivalent to a perfect lottery in top European football.
@@
Keywords: Elo rating; fairness; football; momentum; penalty shootout
@@ 1 Introduction
The penalty shootout is the ultimate tie-breaking rule in (association) football and several other sports to decide which team qualifies in a knockout match. Penalty shootouts usually take place if the result is tied in both regular and extra time. Penalty shootouts directly follow regular time in some tournaments, too. For instance, in the 2025 Leagues Cup, games went straight to a penalty shootout after a tie, and teams were awarded three points for a win in regular time, two points for a win by a penalty shootout, and one point for a loss by a penalty shootout (Leagues Cup,, 2025).
@@
All of these issues have partially been investigated in the existing literature, as revealed by Section 2. However, our research has three innovative aspects. First, we analyse a relatively homogeneous sample provided by matches played in UEFA club competitions without any matches from national cups, where the strengths of the opponents can vary widely. Second, psychological momentum enjoyed by the comeback team is identified in a novel way, according to which team scored the last goal. Third, team strength is measured by Football Club Elo Ratings, a well-established statistical method to create ratings based on past performance (Aldous,, 2017; Gomes de Pinho Zanco et al.,, 2024; van Eetvelde and Ley,, 2019). Previous studies used either the division of the team (Arrondel et al.,, 2019; Krumer,, 2020), or betting odds (Wunderlich et al.,, 2020; Pipke,, 2025) for this purpose. But the first approach is inappropriate in international competitions, while betting odds suffer from several biases (Braun and Kvasnicka,, 2013; Cain et al.,, 2000; Feddersen et al.,, 2017; Winkelmann et al.,, 2021), and may be less persuasive for decision-makers than a measure given by an “objective” mathematical formula.
As we show, the outcomes of the 268 penalty shootouts played in UEFA club competitions between 2000 and 2025 seem to be entirely unpredictable. This contradicts the results of some previous works, which uncover a first-mover advantage (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020) and the positive impact of psychological momentum (Krumer,, 2021). Furthermore, according to our knowledge, all existing studies agree that stronger teams are more likely to win in a shootout (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025), which cannot be observed in UEFA club competitions.
@@ 2 Related literature
The issue of first-mover advantage in football penalty shootouts has received serious attention since the pioneering work of Apesteguia and Palacios-Huerta, (2010). Some studies find evidence for a significant advantage enjoyed by the team kicking the first penalty in each round (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020). These results have inspired formal modelling of the first-mover advantage, as well as several proposals for alternative mechanisms to mitigate or even eliminate this source of unfairness (Palacios-Huerta,, 2012; Echenique,, 2017; Brams and Ismail,, 2018; Vandebroek et al.,, 2018; Del Giudice,, 2019; Anbarcı et al.,, 2021; Csató, 2021a, ; Csató, 2021b, ; Csató and Petróczy,, 2022; Brams et al.,, 2024). The alternating (or ABBA) order was even tried in various tournaments between 2017 and 2018 (Palacios-Huerta,, 2020; Csató, 2021a, ), albeit the IFAB (International Football Association Board), the rule-making body of football, stopped the experiments in 2018 (FIFA,, 2018). It is worth noting that the first-mover advantage seems to be non-existent in tennis tiebreaks, which use the ABBA sequence (Cohen-Zada et al.,, 2018; Da Silva et al.,, 2018).
On the other hand, several papers report no significant difference between the winning probability of the first-mover and the second-mover (Kocher et al.,, 2012; Arrondel et al.,, 2019; Santos,, 2023). According to Kassis et al., (2021), teams whose captains win the coin toss and can choose the shooting order enjoy the advantage. The conflicting results can probably be attributed to small sample sizes (Vandebroek et al.,, 2018). Two recent studies (Vollmer et al.,, 2024; Pipke,, 2025) analysing the highest number of penalty shootouts with respect to first-mover advantage—1759 and 7116, respectively—do not find evidence for the effect of the shooting sequence. However, while the dataset of Pipke, (2025) includes all available shootouts up to the 2023/24 season, this does not mean that first-mover advantage cannot emerge in a more homogeneous subset of matches.
@@
We know only one study on the impact of psychological momentum, although the performance of men is significantly affected by psychological momentum in sports (Cohen-Zada et al.,, 2017). Krumer, (2021) analyses 214 penalty shootouts from two-legged matches in European cups (European Champion Clubs’ Cup/UEFA Champions League, European Cup Winners’ Cup/UEFA Cup Winners’ Cup, UEFA Cup/UEFA Europa League) between 1970 and 2018. Every additional goal scored by the away team in regular time in the second leg increases its probability of winning the penalty shootout by 12.5 percentage points. However, if the home team wins in regular time, it has only a 50% chance to qualify after a shootout.
@@
Stronger football teams have a significantly higher chance of winning a penalty shootout based on 1067 shootouts from the domestic cup competitions of ten European associations (the top five, as well as Belgium, Portugal, Russia, Turkey, Ukraine), the Brazilian cup, the UEFA Champions League, and the UEFA Europa League (Wunderlich et al.,, 2020). However, the effect of team strength on success remains rather small; the winning probability does not exceed 60% even against an extremely weak team. The novelty of Wunderlich et al., (2020) resides in the measure of strength difference, determined by winning probabilities based on pre-match betting odds, adjusted for home advantage. The recent analysis of all available penalty shootouts until the 2023/24 season reinforces that the favourites identified by pre-match odds are more likely to win (Pipke,, 2025, Table A.5). But the favourite team might not be the stronger team due to home advantage (Wunderlich et al.,, 2020).
Brinkschulte et al., (2023) examine the effect of skill on the success of individual penalty kicks. Based on 1574 penalties in major international tournaments, the probability that the goalkeeper saves a penalty that is shot on target decreases if a highly skilled player takes the kick. Skill is a composite measure of player market value, goalkeeper market value, and team quality derived from betting odds.
@@ 3 Data and methodology
Champions League Europa League Conference League Total Qualification 41 (19) 103 (31) 63 (63) 207 (113) Knockout stage 14 (5) 27 (9) 9 (9) 50 (23) Final 6 (0) 5 (3) 0 (0) 11 (3) Total 61 (24) 135 (43) 72 (72) 268 (139) UEFA Conference League was called UEFA Europa Conference League until the 2023/24 season. The row Knockout stage counts only two-legged matches, but not the final, which is played on a neutral field. Numbers in parenthesis show the matches played from the 2020/21 season to 2025.
@@
Figure 1 presents the number of penalty shootouts in each season from 2000/01 to 2025/26 (in the latter case, without the spring of 2026). Besides the natural fluctuation, a dramatic increase can be seen from 2020/21, when most qualification matches were played in one leg instead of two due to COVID-19 restrictions, such that the field of the match was decided by a random draw. In addition, the away goals rule was abolished, and a new competition, the UEFA (Europa) Conference League, started in 2021/22. Therefore, our sample is dominated by recent matches over these 25 years: 139 of the 268 penalty shootouts (51.9%) took place in the last six seasons (Table 1).
@@
The first-mover advantage can be analysed on the basis of all (268) penalty shootouts since this information is always available. On the other hand, some matches were played on a neutral field. In order to examine home advantage, we consider two sets of matches, an extended and a restricted set, where the latter is a strict subset of the former. The extended set does not contain the finals except for the 2011/12 UEFA Champions League final, which was played in München by Bayern München and Chelsea (10 matches are dropped). In addition, one match played by an Israeli and two matches played by Belarusian teams in 2024 and 2025 are disregarded, which leads to 255 matches in the extended set. In the restricted set, 20 matches played behind closed doors in the 2020/21 season, as well as the 2011/12 UEFA Champions League final, are removed.
We identify a comeback team in two different ways. The first adopts the method of Krumer, (2021): the team that scored more goals than its opponent (has a positive goal difference) in the match is the comeback team. Consequently, 121 matches where the second leg game is tied should be ignored. The second approach is novel: the team that scored the last equalising goal, which is responsible for reaching the penalty shootout, is the comeback team. Hence, only 37 shootouts could not be examined.
@@ 4 Results and discussion
As we have seen in Section 3, the number of penalty shootouts has substantially increased from the 2020/21 season. Thus, all binomial tests have also been conducted for the first 20 and the last six seasons, respectively. According to Tables A.3A.6 in the Appendix, the conclusions remain unchanged if we focus on the beginning or the end of the dataset. The only mildly significant value appears for home advantage based on the extended dataset: home teams are less likely to win a penalty shootout than their opponents between 2020 and 2025 if the matches played behind closed doors are taken into account (Table A.5). In these 20 one-legged matches, only seven was won by the home team, which is significantly below 50% at a significance level of 10%. This can be explained by choking under pressure (Harb-Wu and Krumer,, 2019), which was probably much stronger in these matches than usual, despite the absence of spectators, as, in contrast to the standard rule, the randomly drawn home team should not have played away. Even this marginally significant effect disappears when only proper two-legged matches are considered under the restricted definition of home venue (Table A.5).
Table 2: Logistic regression models: the effect of playing at home and team strength
Strength difference $\Delta$ Elo rating $\Delta^{(1)}$ Winning probability $\Delta^{(2)}$ Coefficient -0.121 -0.027 $\beta_{0}$ Standard error 0.123 0.307 $p$-value 0.326 0.930 Coefficient -0.004 -0.194 $\beta_{1}$ Standard error 0.069 0.582 $p$-value 0.959 0.739 Number of observations 266 266 In the model Elo rating, strength difference $\Delta^{(1)}$ is the Elo rating of the home team minus the Elo rating of the away team divided by 100. In the model Elo rating, strength difference $\Delta^{(2)}$ is the winning probability of the home team based on the Elo ratings of the two teams. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
The lack of positive impact is in stark contrast to the findings of the previous literature: all existing papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025) demonstrate a significantly higher winning probability for a “sufficiently” stronger team. Thus, Table 2 investigates the same issue by logistic regressions similar to Wunderlich et al., (2020). Both coefficients are insignificant, that is, neither the field of the match nor the strength of the teams influence success in penalty shootouts, independent of how the strength difference is measured. Furthermore, the estimations of $\beta_{0}$ and $\beta_{1}$ are negative in accordance with Figures 3 and 4, making the implications regarding home advantage and team strength even more robust.
@@ Acknowledgements
We are grateful to Alex Krumer and Fabian Wunderlich for useful comments.
@@ A.2 Supplementary tables
Table A.1: Results of binomial tests: first-mover advantage,
home advantage, momentum from the 2000/01 season to 2025
@@
Table A.2: Results of binomial tests: team strength from the 2000/01 season to 2025
Table A.3: Results of binomial tests: first-mover advantage, home advantage,
momentum from the 2000/01 season to the 2019/20 season
Potential success factor Wins Loses Total Win ratio $p$-value First-mover 72 57 129 55.8% 0.218 Home (extended data) 64 58 122 52.5% 0.651 Home (restricted data) 64 57 121 52.9% 0.586 Comeback (all goals) 40 35 75 53.3% 0.644 Comeback (last goal) 53 55 108 49.1% 0.923 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
@@
from the 2000/01 season to the 2019/20 season
Scenario Favourite wins Favourite loses Total Win ratio $p$-value Elo difference >0>0 60 69 129 46.5% 0.481 Elo difference >25>25 56 64 120 46.7% 0.523 Elo difference >50>50 50 58 108 46.3% 0.501 Elo difference >75>75 42 50 92 45.7% 0.466 Elo difference >100>100 36 40 76 47.4% 0.731 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
Table A.5: Results of binomial tests: first-mover advantage,
home advantage, momentum from the 2020/21 season to 2025
Potential success factor Wins Loses Total Win ratio $p$-value First-mover 71 68 139 51.1% 0.865 Home (extended data) 54 79 133 40.6% 0.037* Home (restricted data) 47 66 113 41.6% 0.090 Comeback (all goals) 32 40 72 44.4% 0.410 Comeback (last goal) 68 55 123 55.3% 0.279 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
Table A.6: Results of binomial tests: team strength from the 2020/21 season to 2025
Scenario Favourite wins Favourite loses Total Win ratio $p$-value Elo difference >0>0 65 72 137 47.4% 0.608 Elo difference >25>25 59 63 122 48.4% 0.786 Elo difference >50>50 51 54 105 48.6% 0.845 Elo difference >75>75 42 45 87 48.3% 0.830 Elo difference >100>100 37 41 78 47.4% 0.734 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.